OpenStudy (anonymous):

if a system has an infinite number of solutions, use set-builder notation to write the solution set. If a system has no solution, state this. 3x + 4y = 5, 2x + y = 1

OpenStudy (anonymous):

3x + 4y = 5,...........................................(1) 2x + y = 1 * by 4 8x + 4y = 4.............................................(2) subtract: (2) - (1) 5x = -1 x = -1/5 plug x = -1/5 in (1) -3/5 + 4y = 5 y = 5.6 / 4 = 1.4 solution is (-0.2,1.4)